Find the final amount in the following retirement​ account, in which the rate of return on the account and the regular contribution change over time.

​$344 per month invested at 4​%,compounded​ monthly, for 6 ​years; then ​$779 per month invested at 6​%, compounded​ monthly, for 6 years.

What is the amount in the account after 12 ​years?
Answer #1

Regular Deposits.

P_1=M_1\left ( \left ( 1+\frac{i_1}{q} \right )^{nq}-1 \right )\left ( \frac{q}{i_1} \right )

After 6 years

Here M_1=\$344,i_1=0.04,q=12,n=6 .

笠,,- 1)(끊) 0.04 P, 344 | | 1 + 0.04 Р, = $27940.562

After 12 years

no no )72 )(da) A-2794 3242(1+0.쓸)72+779(0 +779 1 P2- 27940.562 1+ Ps $107,324.62